OpenStudy (anonymous):

When solving this equation by completing the square, which of the following shows the equation with the square completed? x2 - 4x - 32 = 0 A. (x - 4)^2 = 36 B. (x - 2)^2 = 36 C. (x - 2)^2 = 36

OpenStudy (anonymous):

coefficient of term "x" determines what to add/subtract to complete the square.

OpenStudy (anonymous):

Completing a square requires you to have a square (not surprisingly). So you need the (a-b)^2 form again. However, with the (a-b)^2 rule, you can't have a negative last term. So move the 32 over to the right side of the equation

OpenStudy (anonymous):

GT is right, the coeff of x does determine what extra numbers you need to add/subtract for the whole "complete the square" process. The "formula" to get the last term of the right side of the equation is (|b|/2)^2. B is -4, so putting -4 into the formula yields 4. Make sense?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

Shoot. Where did I lose you?

OpenStudy (anonymous):

All of it Algebra to me is one of the most confusing subjects and it's hard for me to grasp

OpenStudy (anonymous):

haha, yeah we all have our strengths and weaknesses. Are you more of an English/literature person then? Anyway, according to (a-b)^2 = a^2 - 2ab + b^2, and considering that the left side of our equation looks like x^2-4x, wouldn't you say that b You know what? I'm going to try a new approach. Hold on a sec.

OpenStudy (anonymous):

Im more a physical heath and safety with sex ed type of person..

OpenStudy (anonymous):

\[x^{2} - 4x - 32 = 0\]\[x^{2} - 4x = 32 \]\[x^{2} - 4x + [blank] = 32 + [blank]\] To find [blank] use \[({|[coefficient\ of\ x]| \over 2})^{2}\]so [blank] = \[({|-4| \over 2})^{2}=({4 \over 2})^{2} = 2^{2} = 4\] So putting it all together, \[x^{2} - 4x +4 = 32 +4\]

OpenStudy (anonymous):

maybe i should try too lol y=ax^2 +bx +c convert it perfect squared form now (x+b)^2 +C thats the form where b=1/2 * coefficient of x C=c-(1/4*coefficient of b ^2 )

OpenStudy (anonymous):

And those subjects (phys health & sex safety) are much more interesting than the generic math vs reading. Cool! And math once more: You need to factor the x^2 - 4x + 4. To do so (this is the traditional factoring method btw), look for factors of 4 that add up to -4. Factors of 4: 1, 2, 4 AND -1,-2,-4 Of the above factors, only -2 and another -2 will yield a sum of -4. So your equation looks like this now:\[(x-2)^2=32+4=36\]

OpenStudy (anonymous):

Creative I will finish up with u l8ter I got to go to work

OpenStudy (anonymous):

ok, hope you get it later at least :)

OpenStudy (anonymous):

yeah

OpenStudy (callisto):

To start with... completing square is actually making a quadratic equation into the form (x+constant)^2 = a constant. Let's say, it is (x+z)^2 = y Get this part?

OpenStudy (callisto):

, where z and y are constants

OpenStudy (anonymous):

The way you put it Kind of yes

OpenStudy (callisto):

Now, here comes the problem, how to make it into that form? <- your question actually. To begin with, you need to know this identity: \[(a \pm b)^2 = a^2\pm2ab+b^2\] Got it?

OpenStudy (anonymous):

But I am confused where everything go

OpenStudy (callisto):

Oh... first you need to understand that identity... Do you understand that identity ?

OpenStudy (anonymous):

No

OpenStudy (callisto):

:| Okay, let's do it. \[(a+b)^2 = (a+b)(a+b) = a(a+b) +b(a+b)\]\[ = a^2 + ab+ab+b^2 = a^2 + 2ab+b^2\] Got it?

OpenStudy (anonymous):

Oh yeah

OpenStudy (callisto):

So, can you try expanding \[(a-b)^2\]? Let see what you've got for it :)

OpenStudy (anonymous):

ok

OpenStudy (callisto):

@PsYcHoEmOmAtH

OpenStudy (callisto):

@PsYcHoEmOmAtH What have you got so far?

OpenStudy (anonymous):

-4-4^8-32=0

OpenStudy (callisto):

Huh? I was asking you to expand \[(a-b)^2\] ...

OpenStudy (callisto):

As you don't understand the identity, we must do it once first, make sure you understand it, knowing it how it comes

OpenStudy (anonymous):

One minute please my friend is trying to help me out here in class

OpenStudy (callisto):

Hmm.. Sorry to leave you alone .. But I have to sleep now.

OpenStudy (anonymous):

Its ok bye